OpenStudy (anonymous):

There are 4 blue disks and 3 yellow disks in a bag. Jane and Mary play a game where they each have a turn at drawing a disc with replacement. Jane starts. Jane wins when she draws a blue disc and Mary wins when she draws a yellow disc. (a) What's the probability that Jane wins on the first draw? (b) What is the probability that Jane wins in the first 3 draws? (c) What is the probability that Jane wins the game?

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

yes, how do i start?

OpenStudy (anonymous):

first question is easy enough jane starts there are 4 blue disks and 7 total, so the probability she wins on the first draw is \(\frac{4}{7}\) is that part ok?

OpenStudy (anonymous):

yes, what about the second part?

OpenStudy (anonymous):

is the next part \[\frac{4}{7} *\frac{3}{6} *\frac{2}{5} \]?

OpenStudy (anonymous):

ok now we want the probability jane wins in the first 3 draws, which i take to mean either her first draw, her second draw, or her third draw we have already computed the probability she wins on her first now lets compute the probability she wins on her second it is going to take more work don't forget this is with replacement

OpenStudy (anonymous):

how could she win on her second draw? jane draws yellow, \(p=\frac{3}{7}\) then mary draws blue \(p=\frac{4}{7}\) then jane draws blue \(p=\frac{4}{7}\)

OpenStudy (anonymous):

hope it is clear what the logic is and where i got the numbers from in other words it goes jane doesn't win mary doesn't win jane does win that probability is the product of the above numbers, since they all have to occur, and that product is \((\frac{4}{7})^2\times \frac{3}{7}\)

OpenStudy (anonymous):

now you have to do it one more time jane doesn't win mary doesn't win jane doesn't win mary does't win jane DOES win compute each of these probabilities (they are not hard, remember it is with replacement) and then multiply them together use exponential notation and do not actually compute i can write it out for you if you like don't despair, the last question is the easiest

OpenStudy (anonymous):

did you compute the last one? i think that problem is there to get you to see the pattern and have you sum a geometric series for the last problem we can do it a snap way if you like

OpenStudy (anonymous):

is the second one\[\frac{4}{7} \left(\frac{3}{4}*\frac{4}{7}*\frac{4}{7}\right) +\frac{3}{7} *\frac{4}{7} *\frac{3}{7} *\frac{4}{7} *\frac{4}{7} \]? sorry im horrible with probability

OpenStudy (anonymous):

looks good but lets write in exponential notation

OpenStudy (anonymous):

\[\frac{4}{7} + \left(\frac{3}{4}*\frac{4}{7}*\frac{4}{7}\right) +\frac{3}{7} *\frac{4}{7} *\frac{3}{7} *\frac{4}{7} *\frac{4}{7}\] is more like it

OpenStudy (anonymous):

the first \(\frac{4}{7}\) for win on the first try

OpenStudy (anonymous):

clear so far?

OpenStudy (anonymous):

oh yeah woops

OpenStudy (anonymous):

yes, how do i do the next one?

OpenStudy (anonymous):

now lets write it so we see the pattern

OpenStudy (anonymous):

looks like \[\frac{4}{7}+\frac{4}{7}\times \left(\frac{4}{7}\times \frac{3}{7}\right)+\frac{4}{7}\times \left(\frac{4}{7}\times \frac{3}{7}\right)^2\]

OpenStudy (anonymous):

pattern is clear yes? and it will continue you can add this up if you know how to add a geometric series but you can solve the problem a different way as well

OpenStudy (anonymous):

yeah, its a GP. what is the different way that i can solve it?

OpenStudy (anonymous):

we think actually this is rather interesting because it is the probabilistic proof of how to sum a geometric series we are only interested in the probability than jane wins so lets put that as \(J\)

OpenStudy (anonymous):

what could happen? she could win on the first try, and that happens with probability \(\frac{4}{7}\)

OpenStudy (anonymous):

OR she could not win, mary not win. this happens with probability \(\frac{3}{7}\times \frac{4}{7}\) and then it is jane's turn again now her probability of not winning is identical to what it was on her first try

OpenStudy (anonymous):

so we get the equation for \(J\) as \[J=\frac{4}{7}+\frac{4}{7}\times \frac{3}{7}J\]

OpenStudy (anonymous):

i hope the reasoning is clear, why the addition and why the multiplication now if you do the algebra you will see solving for J is identical to summing the geometric series written above

OpenStudy (anonymous):

i got it thanks!

OpenStudy (anonymous):

for the first one you are going to write \[\frac{\frac{4}{7}}{1-\frac{3}{7}\times \frac{4}{7}}\] and you are going to solve the linear equation in exactly the same fashion

OpenStudy (anonymous):

yw